Chapter 9 Applying Excel (B) Saved Requirement 2: Revise the data in your worksheet to reflect the results for the subsequenta. What is the activity variance for revenue? (Indicate the effect of each variance by selecting F for favorable, U for ub. What is the spending variance for the cost of ingredients? (Indicate the effect of each variance by selecting F for favoc. What is spending variance for wages and salaries? (Indicate the effect of each variance by selecting F for favorable, Ud. What is spending variance for total expenses? (Indicate the effect of each variance by selecting F for favorable, U fo
Answer #1

Ans:

a                          
Activity variance for revenue =Actual revenue-Planning budgeted revenue= (1800*16.50)-(1700*16.50)=1650   F  
b                          
Spending variance for the cost of ingredients = Actual  Cost of Ingredients-Budgeted cost for actual meals served= 11140-(1800*6.25)=110   F  
c                          
Spending variance for wages and salaries =Actual  Cost of wages and salaries-Budgeted cost =10640-10400=240   U  
d                          
Actual total expenses =11140+10640+1170+220+2005=27155                   27155      
Flexible budget total expenses = (1800*6.25)+10400+800+(1800*0.2)+2200+600+(1800*0.8)=27050                           27050
Spending variance for total expenses = 27155-27050=105 U